Bond X is a premium bond making semiannual payments. The bond pays a coupon rate of 8 percent, has a YTM of 6 percent, and has 14 years to maturity. Bond Y is a discount bond making semiannual payments. This bond pays a coupon rate of 6 percent, has a YTM of 8 percent, and also has 14 years to maturity. The bonds have a $1,000 par value. 

What is the price of each bond today? If interest rates remain unchanged, what do you expect the price of these bonds to be one year from now? In five years? In 10 years? In 12 years? In 14 years? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)
Answer #1

Calculate the price of the bond as follows:

C 366 Price of the bond Bond X 367 0 $1,187.64 368 1 $1,178.77 $1,137.54 370 101 $1,070.20 $1,037.17 372 $1,000.00 369 Bond Y

Formulas:

366 Price of the bond 367 0 368 1 369 5 370 10 371 12 372 14 Bond X Bond Y =PV(6%/2,(14-A367)*2,-1000*8%/2,-1000) =PV(8%/2,(1